OpenStudy (mysticfury):

What are the x- and y- intercepts for the line y+8 = 4(x−5)? Enter your answers as a whole number The x-intercept is . The y-intercept is

OpenStudy (mysticfury):

i know how to do this normally, you replace the other with a 0 and solve it, but i dont get a whole number when i do that for this problem, im confused :/

OpenStudy (mysticfury):

@3mar @Awolflover1

OpenStudy (mysticfury):

@pooja195 @KamiBug @AloneS

OpenStudy (mysticfury):

"Enter your answers as a whole number" it actually says it does

563blackghost (563blackghost):

Let's first write this in slope-intercept form :) First distribute `4` into the paranthesis. \(\huge\bf{4(x-5) \rightarrow 4 \times x + 4 \times -5}\) What does that equal?

OpenStudy (mysticfury):

yep. 4x - 20

563blackghost (563blackghost):

Now we have... \(\huge\bf{y+8=4x-20}\) we subtract 8 from each side. \(\huge\bf{y=4x-28}\) Got it?

OpenStudy (mysticfury):

yep, i got to that point :P

563blackghost (563blackghost):

X - intercept = `(#,0)` Y - intercept = `(0,#)` To find the `x - intercept` we would plug in 0 as `y` in the slope-intercept equation. \(\huge\bf{0=4x-28}\) Simplify to find x.

OpenStudy (mysticfury):

ah okay i gotcha, so x = 7

OpenStudy (mysticfury):

and y = -28 okay, okay, thats what i got the first time, but i didnt think it was right

563blackghost (563blackghost):

Nicely done ;)

OpenStudy (mysticfury):

thank you kind sir for your help.

563blackghost (563blackghost):

*ma'am* Your welcome :)

OpenStudy (mysticfury):

sorry X3 let me fix it thank you kind madam for your assistance

563blackghost (563blackghost):

Your welcome :)

563blackghost (563blackghost):

All's good the name throws ppl off

OpenStudy (mysticfury):

same with me apparently, everyone thinks im a guy too

OpenStudy (3mar):

Sorry, I was not online! The electricity was off! With @563blackghost, you are in good hands! She did it very well as it should be! Thanks for her!

563blackghost (563blackghost):

Thank you @3mar :)

OpenStudy (3mar):

This is the fact! You are brilliant!